# From Voice to Safety: Language AI Powered Pilot-ATC Communication Understanding for Airport Surface Movement Collision Risk Assessment

### v1

@@ 2.2 Surface Movement Collision Risk Assessment
Aviation systems are highly complex cyber-physical systems, where risk is related to the probability of failure when either a sub-module or a whole system is making inappropriate decisions while exposed to hazards [62]. Risk assessment is defined as the systematic identification and evaluation of the risks posed by possible accident scenarios. It is a tool that supports decision making and therefore, risk management. Risk management is optimization of safety of a system, the verification process and risk acceptance, which support airport operations [63, 7]. Casual methods provide the theoretical framework for aviation risk assessment but are mostly data-driven and rely on data quality. Casual methods such as fault tree analysis [64], event tree analysis [65], and common case analysis [66] are used to quantify the statistical probability of an accident or system component failure, while Bow-Tie analysis [67], Petri nets [68], and Bayesian Belief Networks [69] are employed to assess the anticipated risk associated with changes to the system.
Collision risk between two aircraft seeks to quantify the probability of conflicts between aircraft by evaluating random deviations in position and speed, thereby informing separation standards and safety measures. It is a spatiotemporal problem where each aircraft location is uncertain, and researchers emphasize that ignoring the spatiotemporal uncertainty in motion estimation can lead to inaccurate risk assessments [70]. It is quantified by integrating the joint probability of both agents occupying the same space at the same time, given their uncertain trajectories. On the tactical level, the civil aviation risk assessment model can be divided into midair risk assessment and surface risk assessment.
@@
On the surface level, the collision risk assessment expands to other transportation domains. For instance, [86] propose a probabilistic conflict detection model for vessels that uses probability density functions (PDFs) of predicted positions to quantify the conflict criticality between two trajectories. Each vehicle has a travel time distribution (TTD) over each road segment (link), which is one of the major source of uncertainties. Link travel times are nonnegative and often skewed with a long tail (i.e., occasional heavy delays), making log-normal a natural choice. The route travel time is usually modeled as the sum of link travel times, so its distribution is the convolution of the link-level distributions. [87] analyzed empirical travel times and found that a shifted log-normal distribution fit well for road link travel times. Similar approach is adopted for correlated link delays, where the link times are assumed to be log-normal and an approximate analytic form for the travel time density is derived by applying the Fenton-Wilkinson (FW) approximation [86]. Other location-scale distributions such as the log-logistic [88], log-normal&normal [89] have been used for link travel time modeling.
Airport surface movement refers to aircraft movement on the ground, including taxing, takeoff, landing, and aircraft operations on runways, taxiways, taxilanes, and aprons. [90] introduced the Airport Movement Area Safety System (AMASS) to prevent runway incursions from escalating into collisions. [91] evaluated several aircraft taxi time prediction models, identifying key features that enhance modeling accuracy. At both mesoscopic and macroscopic levels, [92] characterized airport surface flow and developed a cell transmission-based model to replicate its spatial–temporal dynamics. [93] analyzed surface movement data to quantify discrete interactions during taxi operations, conceptualizing these as stages of increasing collision risk, while [94] used ASDE‑X surveillance data to assess the frequency and characteristics of potentially hazardous interactions linked to taxiway geometries and traffic flow constraints. Analysis and classification of runaway incursions based strictly on the risk of scenarios associated with the the state at the start of the incursion have been investigated [95, 96, 97].
@@ 3 Methodology
In this section, we introduce the detailed methodology proposed for language AI powered pilot-ATC communication transcripts understanding for airport surface movement collision risk assessment. We propose the end-to-end pipeline that either goes from speech audios or processed communication transcripts to surface collision risk at potential collision spot on the airport node-graph layout. Figure 2 provides the overview of the workflow, where the training and risk assessment module are shown in detail. For the learning function, we collect and obtain our ATC transcripts dataset and propose the rule-enhanced training pipeline by integrating ATC domain specific rules. The surface movement collision risk assessment model takes the distribution of link travel time parameters as inputs and provide the estimation of conflict probability.
@@ 3.1.1 Data Description
We collect our data for the model training and testing from many resources. LiveATC.net is a website providing live and recorded air traffic control (ATC) audio streams from airports worldwide. It offers both web-based streaming and mobile applications for real-time access to ATC audio, making it a valuable resource for research purpose. We adopt the recently developed, state-of-the-art speech-to-text engine, Whisper, to transcribe the audio file into text format and use them as the training set [101]. Besides that, we also download the public released accident investigation reports from NTSB and use as our validation set. These transcripts are records of communications between pilots and air traffic controllers during specific flights or incidents. They are crucial for accident investigations, as they provide a timeline of communications, instructions, and responses that occurred during the flight. We adopt the Tesseract Optical Character Recognition (OCR) to extract text from these scanned documents [102], which offers advanced multilingual support as well. Lastly, we obtain the communication transcripts published online for the test scenarios of our case studies.
The online NER annotation tool is used for manual annotation of our defined entity types. In this study, we are interested in three entities, (a) callsign: the unique ID of the each flight; (b) aircraft state: the state intention of the aircraft, can be any of holding, taxing, approaching, etc; (c) destination: the intended destination of the aircraft, can be a runway name, or a taxiway name, or a gate. Figure 3 shows the count and share of different entity types in the training, validation, and testing sets. The annotation tool directly exports the json file for NER training.
@@ 3.1.2 Experiment Setup
SpaCy is a fast and efficient open-source NLP toolbox designed for production-ready applications involving text analysis. It’s particularly known for its efficiency, accuracy, and ease of use with NER tasks [103]. In this work, we are specifically interested in, EntityRuler, a flexible SpaCy component that allows us to add their domain specific rules to identify entities that the pretrained models might miss, which we call this process rule override. The benefits of EntityRuler are manifolds, (a) domain-specific customization makes easily adapt models to industry-specific terminology; (b) unlike black-box models, rules are clear and easy to debug, making it transparent and controllable; (c) rules can seamlessly integrate into existing models to boost the overall accuracy without retraining. Figure 4 shows several examples of flight callsign rules in regular expression format. We define these rules by reading the related regulations from the ATC manual defined by FAA Order JO 7110.65 and FAA Order JO 7340.2N.
Embeddings serves as dense vector representations of tokens, words, or characters. Classical word embeddings, such as Word2Vec [104] and GloVe [105], assign a single, context-independent vector to each word, failing to capture the nuances of polysemy and contextual usage. In contrast, contextual embeddings generate dynamic representations for tokens, influenced by the surrounding text, thereby encapsulating both syntactic and semantic properties across diverse contexts. This advancement has significantly enhanced performance in various NLP tasks, including text classification, question answering, and text summarization [106]. Token-level embeddings refer to the fine-grained representation of individual tokens (words or characters) within a sequence. Mathematically, Consider a sequence composed of tokens, each token is mapped to a dense vector representation as,
@@
Tok2Vec is considered as local contextual token embedding, in contrary to Word2Vec [104]. It uses a CNN-based model for context-sensitive embeddings, making it more efficient and suitable for syntactic and semantic tasks.
@@
BERT (Bidirectional Encoder Representations from Transformers) is one of the most well-known embedding model due to its superior performance [108]. Unlike traditional models that processed text in one direction, BERT remarkably introduces bidirectional context understanding by training on masked language modeling (MLM) and next sentence prediction (NSP) tasks.
@@
RoBERTa (Robustly Optimized BERT Pretraining Approach) is the modified version of BERT by removing the NSP task, training on the larger datasets, and using dynamic masking. These improvements result in enhanced performance across multiple benchmarks [111].
@@ 3.1.3 Performance and Complexity Study
The metrics used to quantify the classification accuracy are precision, recall, and micro F1 score. Precision tells about how much we can trust the model positive predictions. Recall informs how well the model is able to find all the actual positive samples. Micro F1 score is used to measure how good the model’s capability to balance between precision and recall.
Embedding Model | Augmented Data | Post Prediction Screening | Precision | Recall | F1 Score |
Tok2Vec (Local-Contexual) | Training w/o ATC Rules | No Override | 0.869 | 0.566 | 0.685 |
Override w/ ATC Rules | 0.800 | 0.684 | 0.738 | ||
Training w/ ATC Rules | No Override | 0.839 | 0.684 | 0.754 | |
Override w/ ATC Rules | 0.841 | 0.763 | 0.800 | ||
BERT (Contexual) | Training w/o ATC Rules | No Override | 0.847 | 0.546 | 0.664 |
Override w/ ATC Rules | 0.781 | 0.658 | 0.714 | ||
Training w/ ATC Rules | No Override | 0.838 | 0.750 | 0.792 | |
Override w/ ATC Rules | 0.839 | 0.822 | 0.831 | ||
RoBERTa (Contextual) | Training w/o ATC Rules | No Override | 0.859 | 0.559 | 0.677 |
Override w/ ATC Rules | 0.805 | 0.678 | 0.736 | ||
Training w/ ATC Rules | No Override | 0.853 | 0.724 | 0.783 | |
Override w/ ATC Rules | 0.855 | 0.816 | 0.835 | ||
mBERT (Multilingual) | Training w/o ATC Rules | No Override | 0.871 | 0.533 | 0.661 |
Override w/ ATC Rules | 0.817 | 0.678 | 0.741 | ||
Training w/ ATC Rules | No Override | 0.869 | 0.783 | 0.824 | |
Override w/ ATC Rules | 0.866 | 0.809 | 0.837 | ||
mRoBERTa (Multilingual) | Training w/o ATC Rules | No Override | 0.872 | 0.539 | 0.667 |
Override w/ ATC Rules | 0.820 | 0.691 | 0.750 | ||
Training w/ ATC Rules | No Override | 0.850 | 0.711 | 0.774 | |
Override w/ ATC Rules | 0.856 | 0.822 | 0.839 | ||
DistilBERT (Distilled) | Training w/o ATC Rules | No Override | 0.856 | 0.546 | 0.667 |
Override w/ ATC Rules | 0.800 | 0.711 | 0.753 | ||
Training w/ ATC Rules | No Override | 0.831 | 0.711 | 0.766 | |
Override w/ ATC Rules | 0.840 | 0.796 | 0.818 | ||
BART (Generative) | Training w/o ATC Rules | No Override | 0.840 | 0.691 | 0.758 |
Override w/ ATC Rules | 0.787 | 0.730 | 0.758 | ||
Training w/ ATC Rules | No Override | 0.845 | 0.822 | 0.833 | |
Override w/ ATC Rules | 0.842 | 0.842 | 0.842 |
Table 1 lists the performance of seven embedding models with four different setups, and Figure 5 better visualizes the same results as histograms.
The results obviously indicate that incorporating ATC rules into the model pipeline, both during training and as a post-prediction override, leads to consistent improvements across various embedding models. In particular, the use of ATC rules tends to boost recall significantly, even though it may sometimes lead to a slight reduction in precision. This trade-off is significant, as evidenced by the increased F1 scores, which balance the contributions of both precision and recall. For instance, the metrics of Tok2Vec moves from training without ATC rules and no override (Precision: 0.869, Recall: 0.566, F1: 0.685) to training with ATC rules and override (Precision: 0.841, Recall: 0.763, F1: 0.800) shows a marked increase in recall and F1.
@@
Figure 6 lists the comparison of time and space complexities of our experiments. The metrics to evaluate the space complexity is the number of parameters in the model, which serves as a proxy for its space complexity and also correlates with its capacity to learn complex representations. For instance, mRoBERTa and BART have parameter counts of 560 and 406 million, respectively, suggesting a significant capacity to capture intricate patterns in the data. This increased capacity, as observed in previous accuracy results, translates into higher performance metrics (i.e., F1 scores) when the models are properly fine-tuned. However, this clearly comes at the cost of increased memory requirements and computational overhead during inference.
In contrast, models like DistilBERT (i.e., 66 million parameters) offer a lightweight alternative. The inference time data also reflects this efficiency, while Tok2Vec models are extremely fast (i.e., 0.02 or 0.03 seconds per inference), they tend to lag in accuracy compared to transformer-based models. DistilBERT strikes a balance, with inference times of approximately 0.6 seconds, making it a viable option when computational resources or real-time constraints are a concern. However, models such as BERT and RoBERTa require 1.5 and 1.1–1.4 seconds, respectively, indicating a trade-off between increased model complexity (and hence better accuracy) and slower processing speeds.

### v2

@@ 2.2 Surface Movement Collision Risk Assessment
Aviation systems are highly complex cyber-physical systems, where risk is related to the probability of failure when either a sub-module or a whole system is making inappropriate decisions while exposed to hazards [63]. Risk assessment is defined as the systematic identification and evaluation of the risks posed by possible accident scenarios. It is a tool that supports decision making and therefore, risk management. Risk management is optimization of safety of a system, the verification process and risk acceptance, which support airport operations [64, 7]. Causal methods provide the theoretical framework for aviation risk assessment but are mostly data-driven and rely on data quality. Casual methods such as fault tree analysis [65], event tree analysis [66], and common case analysis [67] are used to quantify the statistical probability of an accident or system component failure, while Bow-Tie analysis [68], Petri nets [69], and Bayesian Belief Networks [70] are employed to assess the anticipated risk associated with changes to the system.
Collision risk between two aircraft seeks to quantify the probability of conflicts between aircraft by evaluating random deviations in position and speed, thereby informing separation standards and safety measures. It is a spatiotemporal problem where each aircraft location is uncertain, and researchers emphasize that ignoring the spatiotemporal uncertainty in motion estimation can lead to inaccurate risk assessments [71]. Recent work attempts to tackle this challenge by modeling aircraft ground movement as a Markov Decision Process within a digital twin framework, integrating a context-aware speed model trained via imitation learning to improve real-time position estimation from noisy A-SMGCS data [72], and the improved tracking accuracy directly benefits collision risk assessment by providing more reliable temporal predictions for potential conflict nodes. It is quantified by integrating the joint probability of both agents occupying the same space at the same time, given their uncertain trajectories. Overall, on the tactical level, the civil aviation risk assessment model can be divided into midair risk assessment and surface risk assessment.
@@
On the surface level, the collision risk assessment expands to other transportation domains. For instance, [88] propose a probabilistic conflict detection model for vessels that uses probability density functions (PDFs) of predicted positions to quantify the conflict criticality between two trajectories. Each vehicle has a travel time distribution (TTD) over each road segment (link), which is one of the major sources of uncertainties. Link travel times are nonnegative and often skewed with a long tail (i.e., occasional heavy delays), making log-normal a natural choice. The route travel time is usually modeled as the sum of link travel times, so its distribution is the convolution of the link-level distributions. [89] analyzed empirical travel times and found that a shifted log-normal distribution fit well for road link travel times. Similar approach is adopted for correlated link delays, where the link times are assumed to be log-normal and an approximate analytic form for the travel time density is derived by applying the Fenton-Wilkinson (FW) approximation [88]. Other location-scale distributions such as the log-logistic [90], log-normal&normal [91] have been used for link travel time modeling.
Airport surface movement refers to aircraft movement on the ground, including taxiing, takeoff, landing, and aircraft operations on runways, taxiways, taxilanes, and aprons. [92] introduced the Airport Movement Area Safety System (AMASS) to prevent runway incursions from escalating into collisions. [93] evaluated several aircraft taxi time prediction models, identifying key features that enhance modeling accuracy. At both mesoscopic and macroscopic levels, [94] characterized airport surface flow and developed a cell transmission-based model to replicate its spatial–temporal dynamics. [95] analyzed surface movement data to quantify discrete interactions during taxi operations, conceptualizing these as stages of increasing collision risk, while [96] used ASDE‑X surveillance data to assess the frequency and characteristics of potentially hazardous interactions linked to taxiway geometries and traffic flow constraints. Analysis and classification of runaway incursions based strictly on the risk of scenarios associated with the state at the start of the incursion have been investigated [97, 98, 99].
@@ 3 Methodology
In this section, we introduce the detailed methodology proposed for language AI-powered pilot-ATC communication transcripts understanding for airport surface movement collision risk assessment. We propose the end-to-end pipeline that either goes from speech audio or processed communication transcripts to surface collision risk at a potential collision spot on the airport node-graph layout. Figure 2 provides an overview of the workflow, where the training and risk assessment modules are shown in detail. For the learning function, we collect and obtain our ATC transcripts dataset and propose the rule-enhanced training pipeline by integrating ATC domain-specific rules. The surface movement collision risk assessment model takes the distribution of link travel time parameters as inputs and provides an estimation of conflict probability.
@@ 3.1.1 Data Description
We collect our data for the model training and testing from many sources. LiveATC.net is a website providing live and recorded air traffic control (ATC) audio streams from airports worldwide. It offers both web-based streaming and mobile applications for real-time access to ATC audio, making it a valuable resource for research purposes. We adopt the recently developed, state-of-the-art speech-to-text engine, Whisper, to transcribe the audio files into text format and use them as the training set [103]. Besides that, we also download the public released accident investigation reports from NTSB and use them as our validation set. These transcripts are records of communications between pilots and air traffic controllers during specific flights or incidents. They are crucial for accident investigations, as they provide a timeline of communications, instructions, and responses that occurred during the flight. We adopt the Tesseract Optical Character Recognition (OCR) to extract text from these scanned documents [104], which offers advanced multilingual support as well. Lastly, we obtain the communication transcripts published online for the test scenarios of our case studies.
The online NER annotation tool is used for manual annotation of our defined entity types. In this study, we are interested in three types of entities, (a) callsign: the unique ID of each flight; (b) aircraft state: the state intention of the aircraft, which can be any of holding, taxiing, approaching, etc; (c) destination intent: the intended destination of the aircraft, which can be a runway name, or a taxiway name, or a gate. Figure 3 shows the count and share of different entity types in the training, validation, and testing sets. The annotation tool directly exports a json file for NER training.
@@ 3.1.2 Experiment Setup
SpaCy is a fast and efficient open-source NLP toolbox designed for production-ready applications involving text analysis. It’s particularly known for its efficiency, accuracy, and ease of use with NER tasks [105]. In this work, we are specifically interested in EntityRuler, a flexible SpaCy component that allows us to add their domain-specific rules to identify entities that the pretrained models might miss, which we call rule override. The benefits of EntityRuler are manifold, (a) domain-specific customization makes it easy to adapt models to industry-specific terminology; (b) unlike black-box models, rules are clear and easy to debug, making it transparent and controllable; (c) rules can seamlessly integrate into existing models to boost the overall accuracy without retraining. Figure 4 shows several examples of flight callsign rules in regular expression format. We define these rules by reading the related regulations from the ATC manual defined by FAA Order JO 7110.65 and FAA Order JO 7340.2N.
Importantly, our approach remains adaptable beyond the U.S. context. We evaluated multilingual embeddings (mBERT and mRoBERTa) to ensure coverage across non-English and mixed-language transcripts. And while FAA orders served as our initial heuristic base, we recognize that ICAO’s global radiotelephony standards (e.g., ICAO Doc 4444 (PANS‑ATM), ICAO Doc 9432 (the Manual of Radiotelephony), provide widely accepted phraseology that forms a common communication foundation in international operations. ICAO also mandates minimum English language proficiency (Operational Level 4) for pilots and controllers in international airspace to ensure a baseline of clarity across diverse linguistic and accent backgrounds. While regional adaptations do occur, for instance, the FAA’s use of TRACON, ARTCC, ramp, and center versus ICAO terms like FIR, apron, and ACC, the EntityRuler’s regex structure allows effortless extension to incorporate these localized terms. Broadening global rule coverage requires substantive engineering effort and ongoing maintenance of the heuristics class.
Embeddings serve as dense vector representations of tokens, words, or characters. Classical word embeddings, such as Word2Vec [106] and GloVe [107], assign a single, context-independent vector to each word, failing to capture the nuances of polysemy and contextual usage. In contrast, contextual embeddings generate dynamic representations for tokens, influenced by the surrounding text, thereby encapsulating both syntactic and semantic properties across diverse contexts. This advancement has significantly enhanced performance in various NLP tasks, including text classification, question answering, and text summarization [108]. Token-level embeddings refer to the fine-grained representation of individual tokens (words or characters) within a sequence. Mathematically, Consider a sequence composed of tokens, each token is mapped to a dense vector representation as,
@@
Tok2Vec is considered local contextual token embedding, in contrary to Word2Vec [106]. It uses a CNN-based model for context-sensitive embeddings, making it more efficient and suitable for syntactic and semantic tasks.
@@
BERT (Bidirectional Encoder Representations from Transformers) is one of the most well-known embedding models due to its superior performance [110]. Unlike traditional models that processed text in one direction, BERT remarkably introduces bidirectional context understanding by training on masked language modeling (MLM) and next sentence prediction (NSP) tasks.
@@
RoBERTa (Robustly Optimized BERT Pretraining Approach) is the modified version of BERT by removing the NSP task, training on larger datasets, and using dynamic masking. These improvements result in enhanced performance across multiple benchmarks [113].
@@ 3.1.3 NER Performance Comparison
The metrics used to quantify the classification accuracy are precision, recall, and micro F1 score. Precision tells about how much we can trust the model’s positive predictions. Recall informs how well the model is able to find all the actual positive samples. Micro F1 score is used to measure how good the model’s capability is to balance between precision and recall.
| Embedding Model | S1 | S2 | S3 | S4 | ||||||||
| Precision | Recall | microF1 | Precision | Recall | microF1 | Precision | Recall | microF1 | Precision | Recall | microF1 | |
| Tok2Vec (Local-Contextual) | 0.869 | 0.566 | 0.685 | 0.800 | 0.684 | 0.738 | 0.839 | 0.684 | 0.754 | 0.841 | 0.763 | 0.800 |
| BERT (Contextual) | 0.847 | 0.546 | 0.664 | 0.781 | 0.658 | 0.714 | 0.838 | 0.750 | 0.792 | 0.839 | 0.822 | 0.831 |
| RoBERTa (Contextual) | 0.859 | 0.559 | 0.677 | 0.805 | 0.678 | 0.736 | 0.853 | 0.724 | 0.783 | 0.855 | 0.816 | 0.835 |
| mBERT (Multilingual) | 0.871 | 0.533 | 0.661 | 0.817 | 0.678 | 0.741 | 0.869 | 0.783 | 0.824 | 0.866 | 0.809 | 0.837 |
| mRoBERTa (Multilingual) | 0.872 | 0.539 | 0.667 | 0.820 | 0.691 | 0.750 | 0.850 | 0.711 | 0.774 | 0.856 | 0.822 | 0.839 |
| DistilBERT (Distilled) | 0.856 | 0.546 | 0.667 | 0.800 | 0.711 | 0.753 | 0.831 | 0.711 | 0.766 | 0.840 | 0.796 | 0.818 |
| BART (Generative) | 0.840 | 0.691 | 0.758 | 0.787 | 0.730 | 0.758 | 0.845 | 0.822 | 0.833 | 0.842 | 0.842 | 0.842 |
Table 1 lists the performance of seven embedding models with four different setups, and Figure 5 visualizes the same results as histograms. The results obviously indicate that incorporating ATC rules into the model pipeline, both during training and as a post-prediction override, leads to consistent improvements across various embedding models. In particular, the use of ATC rules tends to boost recall significantly, even though it may sometimes lead to a slight reduction in precision. This trade-off is significant, as evidenced by the increased F1 scores, which balance the contributions of both precision and recall. For instance, the metrics of Tok2Vec move from training without ATC rules and no override (Precision: 0.869, Recall: 0.566, F1: 0.685) to training with ATC rules and override (Precision: 0.841, Recall: 0.763, F1: 0.800) show a marked increase in recall and F1.
@@
3.1.4 Embedding Model Complexity Study
Figure 6 lists the comparison of time and space complexities of our experiments. The metrics to evaluate the space complexity are the number of parameters in the model, which serves as a proxy for its space complexity and also correlates with its capacity to learn complex representations. For instance, mRoBERTa and BART have parameter counts of 560 and 406 million, respectively, suggesting a significant capacity to capture intricate patterns in the data. This increased capacity, as observed in previous accuracy results, translates into higher performance metrics (i.e., F1 scores) when the models are properly fine-tuned. However, this clearly comes at the cost of increased memory requirements and computational overhead during inference.
In contrast, models like DistilBERT (i.e., 66 million parameters) offer a lightweight alternative. The inference time data also reflects this efficiency. While Tok2Vec models are extremely fast (i.e., 0.02 or 0.03 seconds per inference), they tend to lag in accuracy compared to transformer-based models. DistilBERT strikes a balance, with inference times of approximately 0.6 seconds, making it a viable option when computational resources or real-time constraints are a concern. However, models such as BERT and RoBERTa require 1.5 and 1.1–1.4 seconds, respectively, indicating a trade-off between increased model complexity (and hence better accuracy) and slower processing speeds.
@@ 3.1.4 Embedding Model Complexity Study
3.1.5 Embedding Sensitivity
| Setup | Mean microF1 | SD across embeddings | MicroF1 Range |
| S1 | 0.683 | 0.032 | 0.097 |
| S2 | 0.741 | 0.014 | 0.044 |
| S3 | 0.789 | 0.027 | 0.079 |
| S4 | 0.829 | 0.014 | 0.042 |
Means/SDs computed across the seven embeddings using microF1; SD is population SD; Range is maximum microF1 minus minimum microF1.
We quantify how NER performance depends on the embedding choice by pooling microF1 across seven encoders under the four pipelines in Table 1 and summarizing cross-embedding variability in Table 2. Without rules (S1: train w/o rules, no override), F1 spans 0.661–0.758 (mean 0.683, SD 0.032, range 0.097), indicating substantial sensitivity to the embedding. Adding only a post-prediction override (S2) already compresses variability (mean 0.741, SD 0.014, range 0.044). Training with rules (S3) improves the mean to 0.789 but retains a wider spread (SD 0.027, range 0.079). The most accurate and stable configuration is S4 (train +rules, +override), where F1 concentrates in 0.800–0.842 (mean 0.829, SD 0.014, range 0.042).
Overall, integrating ATC rules reduces sensitivity to the embedding method by roughly 2.3 (range 0.097 in S1 vs. 0.042 in S4) while increasing mean F1 by 0.146. This pattern is consistent with domain adaptation: ATC-specific heuristics both raise recall and constrain error modes across encoders, thereby mitigating latent-space mismatch (e.g., numerals, callsigns, clipped phraseology, mixed accents). Practically, S4 enables competitive performance even with lighter models (e.g., DistilBERT) for low-latency settings, while larger encoders (e.g., BART, mRoBERTa) yield the highest F1 when compute permits.